# Efficient Long Sequential Low-rank Adaptive Attention for Click-through rate Prediction

### v3

@@ 4. Experiments
In this section, we describe the experimental setup and conduct extensive experiments on both public and industrial dataset to evaluate the performance and the cost of the proposed LERA model.
@@ 4.1. Experiments Setup
Datasets. Our offline model evaluation are conducted on two widely used CTR prediction datasets for research and one real-world industrial dataset.
Amazon dataset [^1] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews. It is derived from user interactions on Amazon’s e-commerce platform(Ni et al., 2019). We use Books collection of reviews, pre-process the data follow (Pi et al., 2020). the recent 10 user behaviors as short-term user sequential features and recent 90 user behaviors as long-term user sequential features.
Taobao dataset [^2] https://tianchi.aliyun.com/dataset/dataDetail?dataId=649. It is based on user behavior logs from Taobao, one of the major e-commerce platform in China and contains 89 million records. we take the click behaviors for each user and sort them according to time in an attempt to construct the behavior sequence.The we use the first 64 behaviors as short-term behavior sequence and the rest of 500 clicked items as long-term sequence features.
Industrial dataset. It is collected from the logs of our real-world industrial e-commerce advertising platform. The dataset encompasses advertising impression logs and corresponding click labels during the period from December 3, 2024, to December 23, 2024.It cotains 2 billion records from 10 million users. The user historical clicked items reach tens of thousands within the past year. We use the recent 64 click behaviors as short-term sequential features and 10000 items as long-term sequential features.
@@
Competitors. We compare the proposed LERA method with mainstream CTR models and various of two-stage models for long-term user behaviors. DIN(W/o long seq.) (Zhou et al., 2018): Only short-term sequences are involved. DIN(Long sequence): We apply the DIN target attention for both short-term and long-term sequences, but this method is not possible to deploy online for the intolerable inference latency. We also implement baseline of two-stage method for long-term user behaviors: SIM (Pi et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), TWIN (Chang et al., 2023).
Parameter Setting. The hidden layers’ sizes are set to a fixed tuple [512, 256, 128]. Each hidden layer follows a Leaky ReLU (Maas et al., 2013) and a BatchNorm Layer (Ioffe, 2015) to accelerate the model training. All methods are trained with Adagrad for optimization. We tuned the learning rate in 0.01, 0.005, 0.001. The batch size is set to 1024 and embedding dimension $d$ is set to 16.
@@ 4.2. Experiment Performance
Overall Perfomance. The evaluation results is shown in Table 2. SIM, EAT, SDIM, TWIN, outperform DIN. That indicates long-term behavior sequence modeling is helpful for more accurate modeling of CTR since it can cover more user’s historical interests. DIN(Long Sequence) has a stable improvement of AUC gain of 0.86%/0.65% compared with the best two-stage method on Amazon and Taobao dataset. We can also see a AUC gain of 0.49% and a GAUC gain of 0.79% in industrial dataset when comparing DIN(Long Sequence) with TWIN. But due to the limitation of online inference latency, we can’t deploy DIN to long sequence for modeling long-term user interests. However, our LREA also outperforms all two-stage methods and it achieves similar results to long sequence target attention, and even exceeds DIN on industrial datasets but its online computation complexity is much less than that of long sequence DIN.
Online A/B Test. We conduct online A/B experiments to evaluate LREA in our online advertising system. The experiment lasts for 12 days and LREA achieves 5.88% CTR and 4.26% RPM gain compared to product model (DIN+SIM).
Hyper-parameters Analysis. We conduct analysis experiments of two important hyper-parameters on our industrial dataset: compression rank $r$ and non-negative loss coefficient $\lambda$. As shown in Table 3. First we fix $\lambda$ to 0.3, tune r in from 32 to 128. The Compression rank of $r=128$ achieves the best performance, and AUC/GAUC drop significantly when $r$ becomes smaller. That indicates that we can’t compression the representation of the original long sequences to a very low rank. We will not test $r$ 128, because although the larger r is, the better performance we get, there should be a appropriate rank for compression to balance cache memory size and model performance since we can’t deploy it online due to model inference latency.
As shown in the second major column of Table 3, we fix $r$ at 128 and adjust the value of $\lambda$ between 0.05 and 0.8. As mentioned earlier, the $\lambda$ parameter controls the nonlinearity and non-negativity in the compression module. If $\lambda$ is too large, it reduces the nonlinearity, resulting in a decline in model performance. Conversely, if $\lambda$ is too small, it skews the loss of the main task of CTR estimation. In our experiments, the model performs best when $\lambda=0.3$.
@@ 5. Conclusion
In this work, we introduced LREA, a novel attention mechanism that enhances CTR prediction in online advertising and recommendation systems. LREA effectively integrates long-term user behavior by using low-rank matrix decomposition and matrix absorption. This mitigates the performance degradation caused by traditional top K selection. Experimental results demonstrate that our model achieves significant improvements over the competitors.

### v4

@@ 4. Experiments
In this section, we describe the experimental setup and conduct extensive experiments on public and industrial datasets to evaluate the performance and cost of the proposed LERA model.
@@ 4.1. Experiments Setup
Datasets. Our offline model evaluations are conducted on two widely used CTR prediction datasets and one industrial dataset.
Amazon dataset [^1] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews. It is derived from user interactions on Amazon’s e-commerce platform(Ni et al., 2019). We use Books collection of reviews, pre-process the data follows (Pi et al., 2020). The recent 10 user behaviors are short-term user sequential features and recent 90 user behaviors are long-term user sequential features.
Taobao dataset [^2] https://tianchi.aliyun.com/dataset/dataDetail?dataId=649. It is based on user behavior logs from Taobao, one of the major e-commerce platform in China and contains 89 million records. We take the click behaviors for each user and sort them according to time to construct the behavior sequence. Then we use the first 64 behaviors as short-term behavior sequential and the rest of 500 clicked items as long-term sequential features.
Industrial dataset. It is collected from the logs of our real-world industrial e-commerce advertising platform. The dataset encompasses advertising impression logs and the corresponding click labels during the period from December 3, 2024, to December 23, 2024. It contains 2 billion records from 10 million users. The user historical clicked items reach tens of thousands within the past year. We use the recent 64 click behaviors as short-term sequential features and 10000 items as long-term sequential features.
@@
Competitors. We compare LERA with the mainstream CTR models as follows: DIN(W/o long seq.) (Zhou et al., 2018): Only short-term sequences are involved. DIN(Long sequence): We apply the DIN target attention for both short and long-term sequences, but this method is not possible to deploy online due to intolerable inference latency. We also implement a two-stage method baseline for long-term user behaviors: SIM (Pi et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), TWIN (Chang et al., 2023).
Parameter Setting. The hidden layer sizes are set to a fixed tuple [512, 256, 128]. Each hidden layer follows a Leaky ReLU (Maas et al., 2013) and a BatchNorm Layer (Ioffe, 2015) to accelerate the model training. All methods are trained with Adagrad for optimization. We tuned the learning rate in 0.01, 0.005, 0.001. The batch size is set to 1024 and embedding dimension $d$ is set to 16.
@@ 4.2. Experiment Performance
Overall Perfomance. The evaluation results is shown in Table 2. SIM, ETA, SDIM, TWIN, outperform DIN. That indicates long-term behavior sequence is helpful for CTR prediction since it can cover more user’s historical interests. DIN(Long Sequence) has a stable improvement of AUC gain of 0.86%/0.65% compared with the best two-stage method on Amazon and Taobao dataset. We can also see a AUC gain of 0.49% and a GAUC gain of 0.79% in industrial dataset when comparing DIN(Long Sequence) with TWIN. But due to the limitation of inference latency, we can’t deploy DIN to long sequences. However, our LREA also outperforms all two-stage methods and it achieves similar results to long sequence target attention, and even exceeds DIN on industrial datasets but its inference complexity is much less than that of DIN(Long sequence).
Hyper-parameters Analysis. We conduct analysis experiments of rank $r$ and coefficient $\lambda$ on our industrial dataset. As shown in Table 3. First we fix $\lambda$ to 0.3, tune r from 32 to 128. The rank of $r=128$ achieves the best performance, and AUC/GAUC drop significantly when $r$ becomes smaller. This indicates that we can’t compression the representation of the original long sequences to a very low rank. We didn’t test $r$ 128, because although the larger r is, the better performance we get, there should be a appropriate rank to balance cache memory size and model performance. Next, we fix $r$ at 128 and adjust the value of $\lambda$ between 0.05 and 0.8. As mentioned earlier, $\lambda$ controls the nonlinearity and non-negativity in the compression module. If $\lambda$ is too large, it reduces the nonlinearity, resulting in a decline in model performance. Conversely, if $\lambda$ is too small, it skews the loss of the main task of CTR estimation. In our experiments, the model performs best when $\lambda=0.3$.
Online A/B Test. We conduct online A/B experiments to evaluate LREA in our online advertising system. The experiment lasts for 12 days and LREA achieves 5.88% CTR and 4.26% RPM gain compared to the product model (DIN+SIM). The inference rt only increases by 3 ms. And In comparison, the inference rt of DIN(Long sequence) will average increases about 30 ms, which causes 15% of user requests to time out.
@@ 5. Conclusion
In this paper, we introduced LREA, a novel attention mechanism that enhances CTR prediction in online advertising and recommendation systems. LREA effectively integrates long-term user behavior by using low-rank matrix decomposition and matrix absorption. This mitigates the performance degradation caused by traditional top K selection. Experimental results demonstrate that our model achieves significant improvements over the competitors.